OpenStudy (anonymous):

I cannot figure out how to do this please help for a medal. just explain please and then tell me if i get the correct answer

OpenStudy (anonymous):

Find the product. Simlify your answer. \[\frac{ -4 }{ 3a } x \frac{ 5b }{ -7 }\]

OpenStudy (fibonaccichick666):

So, what do you think is your first step?

OpenStudy (fibonaccichick666):

or, how would you do \[\frac{3}{4} \times\frac{5}{6}\]?

OpenStudy (fibonaccichick666):

hello @AnonymousBella ?

OpenStudy (anonymous):

for fractions you just multiply across